let g be the function defined by the equation below
g(x) = -x^2+6x

find the following :
g(a+h)
g(-a)
g(sqr(a))
a+g(a)
1/g(a)

Answer by nyc_function(2741)

Just replace x with each given value and solve individually.

I will solve g(a + h) only and you can do the rest.

Let g be the function defined by the equation below

g(x) = -x^2+6x

Let x = (a + h).

g(a + h) = -(a + h)^2 + 6(a + h)

g(a + h) = -(a^2 + 2ah + h^2) + 6a + 6h

g(a + h) = -a^2 - 2ah - h^2 + 6a + 6h

Now, you can do the rest.